OpenStudy (anonymous):

How would you set up and solve, csc(120)?

rishavraj (rishavraj):

@lyndseyjanelle ?????

OpenStudy (anonymous):

Where did the 90 come from?

OpenStudy (lυἶცἶ0210):

you could convert \(\csc(120) \) into \(\Large \frac{1}{sin(120)} \)

OpenStudy (lυἶცἶ0210):

Sin 120 is in the quad II

OpenStudy (anonymous):

So the final answer is sin 120?

OpenStudy (lυἶცἶ0210):

If you mean "solve" as in exact value, then no, you need to find what 1/sin(120) is

rishavraj (rishavraj):

final answer is \[\frac{ 1 }{ \sin 120 }\] plug the value of sin 120 \[\sin 120 = \frac{ \sqrt{3} }{ 2 }\] know tht???

OpenStudy (lυἶცἶ0210):

Way to steal my answer xD But yes, if you know the degrees you will see that 60 is the reference angle for 120.  So draw a 30-60-90 triangle:  Sin=opp/hyp